[Brahms-l] run preparations update

From: Flemming Videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Sat Jan 28 2006 - 17:14:33 EST
At this time the IR access is open, and the local group & CA-D people are preparing the
BRAHMS area for the upcoming 62 GeV run. The following is a brief summary of ongoing efforts.

i)   The DAQ has been resurected; A bad VME module was found, and presently replaced, while leaving one missing for the trigger setup. (FV,RAS)
ii)  The elog has had the nesc. fixed to be able to work for 2006 (HI).
iii) Magnet run and test will be performed the coming week, along with safety checkout
     (CA-D)
iv)  DC checks are performed to identify problem with T3 planes that cannot hold voltage. Chambers are flowing Isobutane to have at nominal voltages. (RAS,DB)
v)   The showerMax detectors are being calibrated using cosmis rays. A small test setup was built for this. The 'left' ZDC SH will be completed this weekend, with the measurements moving to the
right setup (DB,EB,JHL)
vi)  HV has been checked for TPCs and DC in general (RD)

We expect to be part of the early setup phase (end of Feb) to tune the detector, and then again
towards the end of May for the 2 week 62.4 GeV physics run.


--------------------------------------------
Flemming Videbaek
Physics Department 
Bldg 510-D
Brookhaven National Laboratory
Upton, NY11973

phone: 631-344-4106
fax:        631-344-1334
e-mail: videbaek @ bnl.gov


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 28 17:15:03 2006

This archive was generated by hypermail 2.1.8 : Sat Jan 28 2006 - 17:15:18 EST